OpenStudy (anonymous):

Choose the correct answer. Calib received $50 for his birthday that he put in his piggy bank. Each week, he puts 2 more dollars in his bank so that the amount of money in dollars D in his bank is given by D = 2w + 50 where w is the number of weeks. His sister Lara makes a table of reasonable values for w to determine how much money Calib will eventually have. If one of the values for w Lara selects is 14, what would be her corresponding entry for D? A. 264 B. 66 C. 78 D. 57

OpenStudy (anonymous):

um, i am not going to keep helping you, if you keep leaving

OpenStudy (anonymous):

im not leaving?

OpenStudy (anonymous):

ok do you want me to help you with this....

OpenStudy (anonymous):

yes and i have 2 other questions

OpenStudy (anonymous):

and then im done

OpenStudy (anonymous):

i just need help wth these last ones. I allready did three times and keep getting them wrong.

OpenStudy (anonymous):

if number of weeks w = 14, then D = (2 * 14) + 50 = 78 Do you get this???

OpenStudy (anonymous):

I do but see i have dyslexia and the number get all jumble up n my head. And t make me wanna scream!

OpenStudy (anonymous):

I understand how to do it. Its just i cant do it right!

OpenStudy (anonymous):

Choose the correct answer. Calib received $50 for his birthday that he put in his piggy bank. Each week, he puts 2 more dollars in his bank so that the amount of money in dollars D in his bank is given by D = 2w + 50 where w is the number of weeks. His sister Lara makes a table of reasonable values for w to determine how much money Calib will eventually have. If one of the values for w Lara selects is 44, what would be her corresponding entry for D? A. 138 B. 96 C. 77 D. 294

OpenStudy (anonymous):

Can u help me witht hat one too? ^^^^^^^^^^^^^^^^^^^^^^^^^^

OpenStudy (anonymous):

if w = 44, D = (2 * 44) + 50 = 138 Any question you post i will be happy to help you....because i understand the dyslexia, because me cousin and big sister is..so i would be glad to help you...

OpenStudy (anonymous):

OK thankyou very much

OpenStudy (anonymous):

Your welcome....do you need anymore help

OpenStudy (anonymous):

Choose the correct answer. Calib received $50 for his birthday, which he put in the bank. Each week, he puts $2 more in the bank, so the amount of money in his bank is D = 2w + 50, where D is the total amount of money and w is the number of weeks. His sister Lara helped him find how much money he will have after 4, 15, 20, and 36 weeks. The table below shows her results. Find the rate of change using Lara's values for D and w. Explain whether this is a linear model. A. linear: The rate of change is½. B. linear: The rate of change is 2. C. linear: The rate of change varies. D. not linear: The rate of change varies.

OpenStudy (anonymous):

The rate of change is the same as the slope of the formula. Since the equation is given in slope-intercept form: y = mx + b where m = slope, b = y intercept Your rate of change will be 2. Since the equation is a linear equation, it is a linear model so my answer would be.... BBBB

OpenStudy (anonymous):

So its b?

OpenStudy (anonymous):

yes......and make sure you tell me your grade on your quiz.ok!!:)

OpenStudy (anonymous):

I GOT A 100%!! IM SO HAPPY!!! :)

OpenStudy (anonymous):

good, do i geta medal...lol:)

OpenStudy (anonymous):

how do i give you one?

OpenStudy (anonymous):

click best response by my name

OpenStudy (anonymous):

did it do it?

OpenStudy (anonymous):

thanks..

OpenStudy (anonymous):

ohh cool your welcome and thankou!!!

OpenStudy (anonymous):

*thankyou

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

anything else before i go

OpenStudy (anonymous):

Nope im all good! Thankyou very much!!

OpenStudy (anonymous):

ok ur welcome bye..haye wat grade r u in.